OpenStudy (anonymous):

Which of the following is a rational number? square root 15 square root 16, square root 21, square root 22 square root 15 square root 16 square root 21 square root 22

OpenStudy (michele_laino):

hint: \[\Large 16 = {4^2}\] whereas 15, 21, and 22 are not perfect squares

OpenStudy (michele_laino):

so we can write: \[\Large \sqrt {16} = 4 = \frac{4}{1}\]

OpenStudy (michele_laino):

for example we can not find a pair of positive integer, say m and n, such that \[\Large \sqrt {21} = \frac{m}{n}\] with n different from zero

OpenStudy (michele_laino):

integers*

OpenStudy (anonymous):

sorry

OpenStudy (michele_laino):

no worries! :) what is the right option?

OpenStudy (anonymous):

so the answer is 16

OpenStudy (michele_laino):

better is square root of 16

OpenStudy (anonymous):

yes square root 16 is the right answer (:

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

i knew it was the right answer cause thats what i had

OpenStudy (michele_laino):

ok!